Fermi National Laboratory


Accelerator Update

Monday May 3
The day shift began with Operations monitoring store 3464 and the Recycler stash, and with MiniBooNE taking beam.

At 8:18 AM, the Antiproton Source (Pbar) resumed stacking.

All the Main Injector (MI) RF stations tripped off at 11:10 AM due to a "bad status" indication. They all reset and turned on okay.

Pbar experts began an hour and fifteen-minute study period at 1:16 PM. This lowered the stacking rate for the duration of the study.

The Preaccelerator I- Source experts allowed the I- column vacuum to rise to normal atmosphere pressures at 2:13 PM. A Linac expert tuned the Linac to run in this mode.

A Recycler expert began a cooling rate study and a Booster expert began a 1.5-hour cogging study at 4:25 PM.

Operations began sending beam to the Meson MTest experiment at 6:17 PM, then shortly after receiving beam they halted beam to access their enclosure.

A MI expert began a three-hour barrier bucket study at 8:02 PM.

Operations established beam to the Meson MCenter experiment at 8:31 PM.

MTest finished using the beam at 9 PM and asked to be disabled.

Tuesday May 4
The midnight shift began with Operations monitoring store 3464 and the stash, with MI and Recycler conducting studies, and with MiniBooNE and MCenter taking beam.

At 12:19 AM, the interlocks for Meson enclosure MC7 dropped for no apparent reason. Operators had the enclosure interlocked by 1:03 AM.

At 3:42 AM, a Linac RF station (LRF3) tripped off for the first of many times. Operators discovered that the cavity temperature was too high and even after the station had been off for a while hadn't come down. Operators called in Linac technicians. This problem halted all beam.

MCenter asked to have their beamline disabled at 4:13 AM.

Linac techs reported at 7:35 AM that LRF3 had a cavity water problem due to its heat exchanger. Water group techs began cleaning the strainer.

Vacuum techs reported at 8:17 AM that a vacuum pump at Tevatron (TeV) sector C23 was failing.

A Linac tech reported at 9:22 AM that LRF3 was repaired.

Pbar resumed stacking at 9:34 AM.

MiniBooNE resumed taking beam at 9:35 AM.

A Pbar expert began an emittance measurement study period at 9:55 AM, followed by a Debuncher cooling study.

A Booster expert began more cogging studies at 10:13 AM.

TeV experts began end-of-store studies at 11:05 AM.

Operations terminated store 3464 at 12:11 PM.

Operations prepared the Collider Detector at Fermilab (CDF) and D-zero (D0) for access into their collision hall at 12:44 PM. CDF came out of access twenty minutes later.

Pbar completed their studies and Operations attempted to reestablish stacking at 3:12 PM. Unfortunately, a problem with a Debuncher kicker (D:EKIK) delayed the process. An expert found some wrong trigger settings, and then Operations found a problem with the septa timers. Stacking resumed at 3:42 PM.

At 3:58 LRF3 tripped off a few times, but each trip reset okay.

At 5 PM, a Booster kicker (B:MKS04) tripped off and wouldn't reset. Operators bypassed the kicker and called in experts who had the station repaired by 6:30 PM.

At 9:25 PM, a Booster septa (B:DCSEP) tripped off. Operators responded and reported a burned component smell. They called in experts to investigate who replaced a power supply and had the septa operating by 10:53 PM. Stacking resumed.

Operations put the TeV through a dry squeeze at 11:49 PM, before starting shot setup.

Monday May 5
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

At 1:27 AM, Operation reported that a Booster RF station had been tripping off and sparking frequently.

Operations established store 3469 at 2:14 AM with an initial luminosity of approximately 59.5E30. Pbar resumed stacking.

Operations reestablished beam to MCenter at 12:17 AM.

At 2:51 AM, the MI Power Supply (PS) loop tripped due to a MI60S overcurrent indication. It reset okay.

At 3:17 AM, Operations reported that they had been unable to extract a slow spill beam from MI. Operators discovered another tripped power supply (I:H518) that had tripped off with the MI PS loop trip and reset it at 3:46 AM.

At 3:23 AM, the MI PS loop tripped again on the same overcurrent indication. It tripped off two more times and operators called in an expert. At 4:30 AM, it tripped and popped a VCB. This shut off all downstream beam: no stacking.

A vacuum expert called Operations at 7:21 AM to remind them of the failing vacuum pump at C2. The vacuum at C2 was degrading due to this.

Plans for Today
Experts will repair the MI problem and then Operations will resume stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 5/5/2004   email Fermilab